OpenStudy (anonymous):

Need help.. its urgent... !! if a, b are roots of the equation px^2 + qx + r= 0, p is not equal to 0 and p,q,r are in A.P. and 1/a + 1/b = 4 , then find the value of |a-b|

ganeshie8 (ganeshie8):

2q = p+r

ganeshie8 (ganeshie8):

1/a + 1/b = 4 (a+b)/(ab) = 4 -q/r = 4

ganeshie8 (ganeshie8):

then use : (a-b)^2 = (a+b)^2 - 4ab

ganeshie8 (ganeshie8):

once you have the numerical value, take the square root to get the value of |a-b|

OpenStudy (anonymous):

@ganeshie8 can you further explain it.. by the way thnxx... !!

ganeshie8 (ganeshie8):

(a-b)^2 = (a+b)^2 - 4ab = (4ab)^2 - 4ab

ganeshie8 (ganeshie8):

find the value of `ab` and plug it above^

ganeshie8 (ganeshie8):

ab = r/p 2q = p+r -q/r = 4 => r/p = -1/9

ganeshie8 (ganeshie8):

So, ab = -1/9

ganeshie8 (ganeshie8):

(a-b)^2 = (a+b)^2 - 4ab = (4ab)^2 - 4ab = (4/9)^2 + 4/9

ganeshie8 (ganeshie8):

simplify and take the squareroot

ganeshie8 (ganeshie8):

you should get |a-b| = 2sqrt(13)/9